OpenStudy (anonymous):

What is the common ratio of the geometric sequence: 128, -16,2,-1/4 -8 1/8 8 -1/8

Nnesha (nnesha):

: geometric sequence ,to find common ratio \[\large\rm \color{red}{r =\frac{ a_2 }{ a_1} ~~or~~ \frac{ a_4 }{ a_3 }}\] divide any term by previous one :-)

OpenStudy (anonymous):

I got-8 is it right?

OpenStudy (mathmath333):

no

Nnesha (nnesha):

how did you get -8 ?

OpenStudy (anonymous):

I divided 128 by -16

Nnesha (nnesha):

divide any term with the `previous ` one a_2 = 2nd term over a_1 = first term divide 2nd term by first one or 3rd term by 2nd one

OpenStudy (anonymous):

I don't get that

Nnesha (nnesha):

 here is an example \[4,8,16\] i would divide 8 by 4 to get ratio not 4 by 8

OpenStudy (anonymous):

so the answer is 1/8

Nnesha (nnesha):

well 2nd term is -16 and first one is 128 \[\frac{ -16 }{ 128 }= ??\]

OpenStudy (anonymous):

0.125

Nnesha (nnesha):

keep it in fraction focus on sign negative divide by positive = ?

OpenStudy (anonymous):

a negative

Nnesha (nnesha):

right so -16/128 = ?

OpenStudy (anonymous):

-1/8

Nnesha (nnesha):

looks good